# Data Requirement Goal Modeling for Machine Learning Systems

### v1

@@ Abstract.
Background: Machine Learning (ML) has been integrated into various software and systems. Two main components are essential for training an ML model: the training data and the ML algorithm. Given the critical role of data in ML system development, it has become increasingly important to assess the quality of data attributes and ensure that the data meets specific requirements before its utilization. Objective: This work proposes an approach to guide non-experts in identifying data requirements for ML systems using goal modeling. Methodology: In this approach, we first develop the Data Requirement Goal Model (DRGM) by surveying the white literature to identify and categorize the issues and challenges faced by data scientists and requirement engineers working on ML-related projects. An initial DRGM was built to accommodate common tasks that would generalize across projects. Then, based on insights from both white and gray literature, a customization mechanism is built to help adjust the tasks, KPIs, and goals’ importance of different elements within the DRGM. The generated model can aid its users in evaluating different datasets using GRL evaluation strategies. We then validate the approach through two illustrative examples based on real-world projects. Results: The results from the illustrative examples demonstrate that the data requirements identified by the proposed approach align with the requirements of real-world projects, demonstrating the practicality and effectiveness of the proposed framework. Conclusion: The proposed dataset selection customization mechanism and the proposed DRGM are helpful in guiding non-experts in identifying the data requirements for machine learning systems tailored to a specific ML problem. This approach also aids in evaluating different dataset alternatives to choose the optimum dataset for the problem. For future work, we recommend further evaluation of the proposed approach across more ML problem types and contexts, as well as implementing tool support to generate the DRGM based on a chatbot interface.
User stories, large language models, requirements quality, AI requirements, requirements elicitation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Requirements analysis
1. Introduction
@@
Machine learning (ML) is being integrating into various real-life applications, including medical diagnosis, stock market trading, and image recognition applications necessitating an investigation into how ML is addressed during the various stages of software development (r1, ; r2, ; r3, ; r4, ; r7, ; r6, ; r5, ). In contrast to conventional software systems, the behavior of ML-based systems is primarily driven by data and the model constructed from it rather than by predefined functionalities and logic designed by engineers (r1, ; r4, ; r6, ; r2, ). Consequently, in the realm of machine learning (ML), data requirements and properties exert a profound influence on model performance, interpretability, and fairness. Data requirements should play a pivotal role in guiding feature selection, preprocessing, and ensuring that the data is effective and safe for learning and prediction.
@@
Understanding data properties, such as distribution and correlations, is essential for model interpretabilityand significantly impacts model evaluation and validation. Addressing bias and fairness in ML models depends on a comprehensive understanding of data properties to effectively mitigate biases. One example that highlights the importance of data properties to building an effective ML system, is a systemic literature review investigating the clinical viability of machine learning models developed to detect and diagnose COVID-19 from chest x-rays in studies published in 2020. This review found that none of the included models demonstrated potential clinical utility (Roberts2021, ). This was due to methodological flaws or underlying biases, related related to the datasets issues, including the use of public datasets where the integrity of the data is questioned, challenges with training data size, balancedness, and the presence of the control group. Another example is the case of the Amazon INC recruiting ML system. Although the gender of the applicant was protected, the models were able to recognize patterns in women’s writing style, which led to penalizing female candidates and, consequently, presenting male candidates as more viable (r30, ).
@@
Although many requirements engineering (RE) methods exist to capture and analyze functional and non-functional requirements of systems of with varying levels of complexity and application domains, with the advancement in technologies and data-driven methods, collecting traditional requirements is no longer sufficient to ensure high quality and sound outputs of systems. It is argued that there is a need to enhance existing RE activities or propose new methods to adapt to the inductive nature of ML requirements, deal with the continuously changing requirements in ML models, and manage various challenges such as performance drift and ethics (r8, ; r4, ; r3, ; r9, ). To ensure robustness, reliability, and fairness in ML systems, new categories of requirements—particularly those related to data and ethics—should be systematically collected and analyzed. Although recent research has addressed the use of goal modeling in ML projects (Gike2022, ; 9218187, ; 6290700, ; Neace2017, ; DIMITRAKOPOULOS201928, ; city23738, ), these studies have not addressed capturing data requirements and analyzing alternative datasets. Therefore, the aim of this study is to propose an approach for identifying and capturing data requirements that must be met prior to using a dataset in the development of ML systems.
@@
To achieve our goal, in this paper, we conducted a literature review to identify the key data requirements necessary for building the Data Requirements Goal Model (DRGM). We then developed a customization mechanism for the DRGM to adapt to various ML problems and contexts. This customization mechanism was designed based on insights from both gray and white literature, as well as expert input, and includes two UML activity diagrams that assist requirements engineers in the customization process. The resulting DRGM identifies essential data requirements for ML systems and supports requirements engineers in evaluating dataset alternatives. The effectiveness of the proposed DRGM and customization mechanism was assessed through two case studies in different contexts. This work is expected to benefit requirements engineers, especially those with limited ML expertise, by aiding in data planning and ensuring that datasets meet critical requirements. The key contributions of this work are as follows:
  1. (1)
@@
This paper is organized as follows. Section 2 presents the literature review, addressing the challenges encountered in developing data requirements for ML systems. Section 3 details the methodology of this study, including the construction of the DRGM and the customization mechanism, along with illustrative examples that demonstrate their effectiveness. Finally, Section 5 presents the conclusion, discusses study limitations, and offers suggestions for future work.
2. Related Work
@@
Recent literature reviews emphasized the importance of requirements processes for eliciting requirements specific to ML systems (habiba2024mature, ; Khloodslr, ). Belani et al. proposed the RE4AI taxonomy, which mapped ML system challenges related to data, models, and systems to various RE processes from a requirements engineering perspective (r1, ). Cerqueira et al. (Cerqueira2022guide, ) presented the RE4AI framework, which integrated ethical considerations into the requirements elicitation process for AI systems. Their study emphasized the the lack of adequate training in AI ethics among software development teams and highlighted the need for greater focus on ethics throughout the software development phases. Silva et al. (de2022requirements, ) focused on technical aspects of requirements elicitation for AI, outlining tools and techniques essential for capturing AI-specific data requirements. Their findings stressed the importance of adaptability in requirements models to accommodate diverse data attributes. Vogelsang and Borg (r2, ) and Horkoff et al. (r8, ) agreed that ML-based systems had shifted the development paradigm from coding to training, suggesting that RE practices needed to evolve to address this shift (r2, ).
To address the importance of capturing AI-specific requirements, recent work (Gike2022, ; 9218187, ; 6290700, ; Neace2017, ; DIMITRAKOPOULOS201928, ; city23738, ) highlighted the role of goal-oriented modeling in refining requirements elicitation and analysis for ML systems and in addressing the unique challenges posed by ML requirements. These studies, however, revealed notable differences in focus areas and methodologies. For instance, $i^{*}$ (yu2010social, ) was applied to elicit requirements and model concepts related to AI applications for individuals with dementia (city23738, ) and for rehabilitation care (Gike2022, ). FLAGS (10.1109/RE.2010.25, ) was employed in (6290700, ) to elicit requirements for a surgical robotic assistant, with the resulting goal model subsequently converted to UML to support the development process.
@@
As for analysis, Ishikawa and Matsuno (r5, ) presented GORE-MLOps to capture the uncertainty and unpredictability inherent in ML systems during implementation. GORE-MLOps modeled different scenarios to meet top-level goals by extending GRL to include three states: feasibility unknown, feasibility validated, and feasibility invalidated. Initially, each goal was assigned the status of feasibility unknown, which was then updated to feasibility validated or feasibility invalidated based on experimental or implementation results. The study introduced the terms proved, denied, or unproved at the leaves of the goal model tree, indicating whether the initial contribution scores assigned to goals were confirmed by experiments. The method demonstrated its effectiveness through an illustrative example.
@@
One of the most recent works in this area was presented by Barrera et al. (10.1016/j.csi.2023.103806, ). The work extended $i^{*}$ to capture ML requirements and presented a metamodel that included ML concepts such as MLGoals, MLTask, Indicator, Dataset, and MLQualityAspects. This metamodel was constructed by the authors and to used through a requirements questionnaire and in collaboration with an ML expert to ensure coverage of all relevant aspects of the ML solution. The metamodel was validated through two use cases: one in an industrial context and another in healthcare. In both cases, the metamodel significantly contributed to exploring project goals, identifying key non-functional requirements, specifying required dataset attributes, and recommending algorithms within a defined subset.
Despite the range of work that addressed goal-oriented modeling for eliciting and analyzing requirements for ML-based systems, there is a lack of studies focusing on data requirements in terms of elicitation, evaluation of alternatives, or trade-offs. Thus, this work aimed to bridge this gap by providing an approach for eliciting data requirements and analyzing alternative datasets.
3. Research Methodology
@@
Figure 1. Overview of the research methodology
3.1. Data Requirements Challenges
We first identify the key data challenges that requirements engineers may encounter when working on ML systems through surveying the related work . These challenges were then mapped into four primary categories representing data properties, as outlined by Amershi et al. (r11, ): data quantity, data quality, data management, and data ethics. Table 1 provides a summary of the identified challenges, along with their descriptions and corresponding categories.
Table 1. Data Requirements Challenges
@@ 3.1. Data Requirements Challenges
Data Quantity Data availability the existence of the data  (r11, ; r1, ; r13, ; r8, ; 9796389, )
Data accessibility the data can be reached  (r11, )
Data Quality Data accuracy data values are correct  (r11, ; r6, ; r2, )
@@
Data freshness concerns the latency when receiving the data, and how long it takes to process incoming data  (r11, ; 9796389, )
@@
Data representativeness assumes that the dataset covers the investigated population by having sufficient and similar distribution in the training and testing datasets  (r6, ; 9796389, )
Data balancedness it is a concern that the number of samples per category should be balanced  (r1, ; r6, ; 9796389, )
Data completeness means that data covers all possible values of the context  (r6, ; r2, )
Data consistency refers that all data in the dataset should be in the same format and representation  (r2, )
@@
Data Management Data logging collecting and saving data over time while keeping track of its metadata  (r11, ; r3, )
Data security refers to the confidentiality and integrity of the data  (r8, ; r4, )
@@
Data Ethics Data discrimination presence of protected attributes or their proxies or the data is biased towards a certain group  (r2, )
Data legality constraints on obtaining and using the dataset  (r2, ; 9796389, )
@@
Data privacy data should not be shared or used for other purposes, especially personal data  (r1, ; 9796389, ; r8, ; r4, ; r7, )
@@
Data safety protecting data from risk or uncertainty and preserving sensitive data  (r1, ; r4, ; r7, )
3.2. Data requirement goal model (DRGM)
@@
The Data Requirements Goal Model (DRGM), shown in Fig. 2, was constructed by mapping the data requirements challenges listed in Table 1 to GRL elements. We used soft goals to capture the essential data properties and qualities required when preparing datasets for ML systems. We those soft goals were to satisfy the data actor, which will be the basis for our evaluation. Additionally, tasks proposing solutions for these data requirement challenges were gathered from the literature and experts in the ML field. The relationships between the DRGM elements are illustrated using contribution and decomposition links in Fig. 3111Higher resolution images are in: https://anonymous.4open.science/r/DRGM-F034/. Initial importance values were assigned to the DRGM elements, and these values will be reassigned using the customization mechanism based on the ML problem type and context.
Figure 2. Initial data requirements goal model.
Figure 3. Operationalization of the data requirements goal model.
@@
Data Quantity is composed of two subgoals: Data Availability and Data Accessibility. The Data Availability goal was further decomposed into a subgoal for identifying the data source and a KPI element for data size, with a ”make” contribution to ensure the goal’s satisfaction. Having a large dataset when training a model from scratch is essential. A small dataset may amplify outliers and fail to capture the full variance in the sample space (r2, ).
The data source also impacts data quality goals. For example, using public data may compromise Data Accuracy, as public datasets can contain labels from non-experts, potentially leading to incorrect data (r2, ). Conversely, collecting data in-house can enhance data accuracy and consistency (r2, ; r11, ). In certain contexts, obtaining data from an authoritative source is also crucial (r11, ).
Data Quality consists of four subgoals: Data Consistency, Data Completeness, Data Balancedness, and Data Freshness. Data Consistency can be achieved through data preprocessing to refine the dataset (r2, ) and by removing redundant data. For Data Balancedness, users can either check the balance of the dataset and assign a percentage to the contributing KPI or treat the dataset for balance and then assign the percentage to the KPI.
@@
Data Completeness is achieved when the data covers the entire range of possible values relevant to the problem. It is further decomposed into the subgoal of Data Accuracy, as inaccurate values can render the data incomplete, and a task to ’Treat Missing Data,’ which can involve interpolation or other techniques. For Data Representativeness, the model should be built and tested on data that represents the target phenomena with a similar data distribution. It is essential to ensure that the dataset includes all target classes or value ranges. Additionally, context-based representativeness—such as spatial, seasonal, demographic, or other domain-specific factors—must be ensured, as a lack of representativeness could lead to data discrimination (r2, ; r6, ).
Data Freshness is the final component of data quality. Its importance varies depending on the context, which will be discussed later. Two tasks contribute to Data Freshness: handling incoming data correctly and obtaining data from sustainable sources (r11, ).
Data Management comprises two subgoals: Data Security and Data Logging. Data Logging is crucial in many ML applications where a stream of data is available to improve the model. This requires a method to handle and log incoming data so that new data is not mixed with already trained data (r11, ). Data Security is equally important to protect and preserve the integrity of the data.
Data Ethics consists of four subgoals: Data Legality, Data Privacy, Data Safety, and Data Free from Discrimination. Data Legality is critical in specific contexts, such as finance, where certain domains have strict legal requirements (r2, ). Consequently, a task to confirm compliance with context-specific legal constraints was added to the Data Legality soft goal.
Ensuring data is free from discrimination is vital, as biases against specific groups can propagate through the model, causing fairness issues (10.1145/3468264.3468536, ). It is therefore essential to identify features that could lead to discrimination and safeguard them, such as gender (r2, ).
3.3. Customization mechanism of the DRGM

### v2

# DRGM: Data Requirement Goal Modeling for ML-Based Systems

@@ Abstract
Given the critical role of data in Machine Learning (ML)-based system development, it has become increasingly important to assess the quality of data attributes and ensure that the data meets specific requirements before its utilization. This work proposes an approach to guide non-experts in identifying data requirements for ML systems using goal modeling. In this approach, we first develop the Data Requirement Goal Model (DRGM) by surveying the scientific literature to identify and categorize the issues and challenges faced by data scientists and requirement engineers working on ML-related projects. An initial DRGM was built to accommodate common tasks that would generalize across projects. Then, a customization mechanism is built to help adjust the tasks, KPIs, and goals’ importance of different elements within the DRGM. The generated model can aid its users in evaluating different datasets using Goal-oriented Requirement Language (GRL) evaluation strategies. We then validate the approach through two illustrative examples based on real-world projects. The results demonstrate that the data requirements identified by the proposed approach align with the requirements of real-world projects, showing the practicality and effectiveness of the proposed framework. For future work, we recommend further evaluation of the proposed approach across more ML problem types and contexts, as well as implementing tool support for generating the DRGM via a chatbot interface.
Index Terms:
Goal model, data requirements, data quality
I Introduction
@@
Machine Learning (ML) is being integrated into various real-life applications, including medical diagnosis, stock market trading, and image recognition applications necessitating an investigation into how ML is addressed during the various stages of software development [7, 47, 33, 28, 30, 38, 27]. In contrast to conventional software systems, the behavior of ML-based systems is primarily driven by data and the model constructed from it rather than by predefined functionalities and logic designed by engineers [7, 28, 38, 47]. Consequently, in the realm of ML, data requirements and properties exert a profound influence on model performance, interpretability, and fairness. Data requirements should play a pivotal role in guiding feature selection, preprocessing, and ensuring that the data is effective and safe for learning and prediction.
@@
Understanding data properties, such as distribution and correlations, is essential for model interpretability and significantly impacts model evaluation and validation. Addressing bias and fairness in ML models depends on a comprehensive understanding of data properties to effectively mitigate biases. One example that highlights the importance of data properties to building an effective ML system, is a systemic literature review investigating the clinical viability of machine learning models developed to detect and diagnose COVID-19 from chest x-rays in studies published in 2020. This review found that none of the included models demonstrated potential clinical utility [41]. This was due to methodological flaws or underlying biases, related to the datasets issues, including the use of public datasets where the integrity of the data is questioned, challenges with training data size, balancedness, and the presence of the control group. Another example is the case of the Amazon INC recruiting ML system. Although the gender of the applicant was protected, the models were able to recognize patterns in women’s writing style, which led to penalizing female candidates and, consequently, presenting male candidates as more viable [15].
@@
Although many requirements engineering (RE) methods exist to capture and analyze functional and non-functional requirements of systems with varying levels of complexity and application domains, with the advancement in technologies and data-driven methods, collecting traditional requirements is no longer sufficient to ensure high quality and sound outputs of systems. It is argued that there is a need to enhance existing RE activities or propose new methods to adapt to the inductive nature of ML requirements, deal with the continuously changing requirements in ML models, and manage various challenges such as performance drift and ethics [25, 28, 33, 42]. To ensure robustness, reliability, and fairness in ML systems, new categories of requirements—particularly those related to data and ethics—should be systematically collected and analyzed. Although recent research has addressed the use of goal modeling in ML projects [21, 9, 39, 18, 32], these studies have not addressed capturing data requirements and analyzing alternative datasets. Data requirements define the necessary properties, constraints, and quality attributes of training data of an ML model based on the problem to be solved. These requirements guide the elicitation, specification, and verification of data needed to support the development and deployment of ML-based systems [17]. Therefore, the aim of this study is to propose an approach for identifying and capturing data requirements that must be met prior to using a dataset in the development of ML systems.
To achieve our goal, we adopted Goal-oriented Requirement Language (GRL) because of its effectiveness to represent objectives, constraints, and trade-offs capability. GRL would allow the requirements engineers along with the data scientists to link the data requirements to higher-level system goals hence ensuring that critical data properties are systematically identified and analyzed in the context of the ML-based system’s success criteria. Moreover, GRL was selected because it enables the evaluation of alternative datasets using formal evaluation strategies to see which dataset better satisfies the defined ML-based system’s goals.
@@
Therefore, we conducted a literature review to identify the key data requirements necessary for building the Data Requirements Goal Model (DRGM). We then developed a customization mechanism for the DRGM to adapt to various ML problems and contexts. This customization mechanism was designed based on insights from both gray literature and scientific literature, as well as expert input, and includes two UML activity diagrams that assist requirements engineers in the customization process. The resulting DRGM identifies essential data requirements for ML systems and supports requirements engineers in evaluating dataset alternatives. The effectiveness of the proposed DRGM and customization mechanism was assessed through two case studies in different contexts. This work is expected to benefit requirements engineers, especially those with limited ML expertise, by aiding in data planning and ensuring that datasets meet critical requirements. The key contributions of this work are as follows:
  1. 1.
@@
This paper is organized as follows. Section II presents the literature review, addressing the challenges encountered in developing data requirements for ML systems. Section III details the methodology of this study, including the construction of the DRGM and the customization mechanism, along with illustrative examples that demonstrate their effectiveness. Finally, Section V presents the conclusion, discusses study limitations, and offers suggestions for future work.
II Related Work
@@
Recent literature reviews emphasized the importance of requirements processes for eliciting requirements specific to ML systems [23, 2]. Belani et al. proposed the RE4AI taxonomy, which mapped ML system challenges related to data, models, and systems to various RE processes from a requirements engineering perspective [7]. Cerqueira et al. [45] presented the RE4AI framework, which integrated ethical considerations into the requirements elicitation process for AI systems. Their study emphasized the lack of adequate training in AI ethics among software development teams and highlighted the need for greater focus on ethics throughout the software development phases. Silva et al. [16] focused on technical aspects of requirements elicitation for AI, outlining tools and techniques essential for capturing AI-specific data requirements. Their findings stressed the importance of adaptability in requirements models to accommodate diverse data attributes. Vogelsang and Borg [47] and Horkoff et al. [25] agreed that ML-based systems had shifted the development paradigm from coding to training, suggesting that RE practices needed to evolve to address this shift [47].
@@
To address the importance of capturing AI-specific requirements, recent work [21, 27, 9, 39, 18, 32] highlighted the role of goal-oriented modeling in refining requirements elicitation and analysis for ML systems and in addressing the unique challenges posed by ML requirements. These studies, however, revealed notable differences in focus areas and methodologies. For instance, $i^{*}$ [51] was applied to elicit requirements and model concepts related to AI applications for individuals with dementia [32] and for rehabilitation care [21]. FLAGS [4] was employed in [9] to elicit requirements for a surgical robotic assistant, with the resulting goal model subsequently converted to UML to support the development process. As for analysis, Ishikawa and Matsuno [27] presented GORE-MLOps to capture the uncertainty and unpredictability inherent in ML systems during implementation. GORE-MLOps modeled different scenarios to meet top-level goals by extending GRL to include three states: feasibility unknown, feasibility validated, and feasibility invalidated. Initially, each goal was assigned the status of feasibility unknown, which was then updated to feasibility validated or feasibility invalidated based on experimental or implementation results. The study introduced the terms proved, denied, or unproved at the leaves of the goal model tree, indicating whether the initial contribution scores assigned to goals were confirmed by experiments. The method demonstrated its effectiveness through an illustrative example.
@@
One of the most recent works in this area was presented by Barrera et al. [5]. The work extended $i^{*}$ to capture ML requirements and presented a metamodel that included ML concepts such as MLGoals, MLTask, Indicator, Dataset, and MLQualityAspects. This metamodel was constructed by the authors and used through a requirements questionnaire and in collaboration with an ML expert to ensure coverage of all relevant aspects of the ML solution. The metamodel was validated through two use cases: one in an industrial context and another in healthcare. In both cases, the metamodel significantly contributed to exploring project goals, identifying key non-functional requirements, specifying required dataset attributes, and recommending algorithms within a defined subset. Despite the range of work that addressed goal-oriented modeling for eliciting and analyzing requirements for ML-based systems, there is a lack of studies focusing on data requirements in terms of elicitation, evaluation of alternatives, or trade-offs. Thus, this work aimed to bridge this gap by providing an approach for eliciting data requirements and analyzing alternative datasets.
III Research Methodology
@@
Figure 1: Overview of the Research Methodology
III-A Data Requirements Challenges
We first identify the key data challenges that requirements engineers may encounter when working on ML systems through a conventional literature review. These challenges were then mapped into four primary categories representing data properties, as outlined by Amershi et al. [3]: data quantity, data quality, data management, and data ethics. Table I provides a summary of the identified challenges, along with their descriptions and corresponding categories.
TABLE I: Data Requirements Challenges
@@ III-A Data Requirements Challenges
Data Quantity Data availability The existence of the data [3, 7, 14, 25, 44]
Data accessibility The data can be reached [3]
Data Quality Data accuracy Data values are correct [3, 38, 47]
@@
Data freshness Concerns the latency when receiving the data, and how long it takes to process incoming data [3, 44]
@@
Data representativeness Assumes that the dataset covers the investigated population by having sufficient and similar distribution in the training and testing datasets [38, 44]
Data balancedness The number of samples per category should be balanced [7, 38, 44]
Data completeness Data covers all possible values of the context [38, 47]
Data consistency All data in the dataset should be in the same format and representation [47]
@@
Data Management Data logging Collecting and saving data over time while keeping track of its metadata [3, 33]
Data security Confidentiality and integrity of the data [25, 28]
@@
Data Ethics Data discrimination Presence of protected attributes (or proxies) or data biased toward a certain group [47]
Data legality Constraints on obtaining and using the dataset [47, 44]
@@
Data privacy Data should not be shared or used for other purposes, especially personal data [7, 44, 25, 28, 30]
@@
Data safety Protecting data from risk or uncertainty and preserving sensitive data [7, 28, 30]
III-B Data Requirement Goal Model (DRGM)
@@
The Data Requirements Goal Model (DRGM), shown in Fig. 2, was constructed by mapping the data requirements challenges listed in Table I to GRL elements. We used soft goals to capture the essential data properties and qualities required when preparing datasets for ML systems. Those soft goals are to satisfy the data requirements for a ML-based system, which are encapsulated in the data actor that will be the basis for our evaluation. Additionally, suggested methods to handle these data requirements challenges were captured from gray and scientific literature. Authors were reporting them as part of the preprocessing steps or as recommendations to enhance a ML model’s performance. The relationships between the DRGM elements are illustrated using contribution and decomposition links in Fig. 3111Higher resolution images are in: https://github.com/asmayamani/DRGM. Initial importance values were assigned to the DRGM elements subjectively by the authors after surveying the literature, and these values will be reassigned using the customization mechanism based on the ML problem type and context.
Figure 2: Original Data Requirements Goal Model
Figure 3: Operationalization of the Data Requirements Goal Model
@@
Data Quantity is composed of two subgoals: Data Availability and Data Accessibility. The Data Availability goal was further decomposed into a subgoal for identifying the data source and a KPI element for data size, with a ’make’ contribution to ensure the goal’s satisfaction. Having a large dataset when training a model from scratch is essential. A small dataset may amplify outliers and fail to capture the full variance in the sample space [47]. The Data Accessibility is the second decomposition of the Data Availability as the availability of the data does not ensure that it is usable, reachable, or understandable. The data source also impacts data quality goals. For example, using public data may compromise Data Accuracy, as public datasets can contain labels from non-experts, potentially leading to incorrect data [47]. Conversely, collecting data in-house can enhance data accuracy and consistency [47, 3]. In certain contexts, obtaining data from an authoritative source is also crucial [3]. Therefore, it is a critical task to ensure that data is accurate.
@@
Data Quality consists of four subgoals: Data Consistency, Data Completeness, Data Balancedness, and Data Freshness. Data Consistency can be achieved through data preprocessing such as resolving inconsistencies to refine the dataset [47] and by removing redundant data. For Data Balancedness, users must first evaluate whether the dataset is balanced with respect to the contributing KPI. If imbalances are found, the dataset should be treated for balancedness before assigning the KPI percentage. Data Completeness is achieved when the data represents the entire range of possible values relevant to the problem. The subgoal Data Accuracy also contributes to it as inaccurate values can render the data incomplete, and a task to ’Treat Missing Data’, which can involve interpolation or other techniques. For Data Representativeness, the model should be built and tested on data that represents the target phenomena with a similar data distribution. It is essential to ensure that the dataset includes all target classes or value ranges. Additionally, context-based representativeness such as spatial, seasonal, or other domain-specific factors must be ensured, as a lack of representativeness could lead to data discrimination [47, 38]. Data Freshness is the final component of data quality. Its importance varies depending on the context, which will be discussed later. Two tasks contribute to Data Freshness: handling incoming data correctly and obtaining data from sustainable sources [3].
Data Management comprises two subgoals: Data Security and Data Logging. Data Logging is crucial in many ML applications where a stream of data is available to improve the model. Therefore, this requires a method to handle and log incoming data with versioning the data so that new data is not mixed with already trained data [3]. Data Security is equally important to protect and preserve the confidentiality, integrity, and availability of data. Data Ethics consists of four subgoals: Data Legality, Data Privacy, Data Safety, and Data Free from Discrimination. Data Legality is critical in specific contexts, such as finance, where certain domains have strict legal requirements [47]. Consequently, a task to confirm compliance with context-specific legal constraints was added to the Data Legality soft goal. Ensuring data is free from discrimination is vital, as biases against specific groups can propagate through the model, causing fairness issues [8]. It is therefore essential to identify features that could lead to discrimination and protect them, such as gender [47]. Nonetheless, Data Privacy is critical data ethics concern that ensure individuals’ cannot be identified. Moreover, protecting them from corruption and ensuring their validity are ethical aspects represented in the Data Safety.
III-C Customization Mechanism of the DRGM